If the Ka of a monoprotic weak acid is 8.8 × 10-6, what is the pH of a 0.36 M solution of this acid?
Answer #1

The corresponding equation will be:

HA + H2O <--> A^-1 + H3O^+1


Ka = [A-][H3O+] / [HA] = 8.8x10^-6

At equilibrium, Let X = [H30+] and [A-] = X and [HA] = 0.36 - X

X^2 / (0.36 - X) = 8.8x10^-6

Since it is a weak acid, so assume X << 0.36 therefore

0.36-X = 0.36


X^2/0.36 = 8.8x10^-6
X^2 = 3.168 x10^-6
X = 1.78 x10^-3 = [H+]

pH = -log[H+]

pH = - log (1.78 x 10^-3)

pH = 2.74